# On Discovering Algorithms for Adversarial Imitation Learning

### v1

@@ 6.1 Evolutionary Search
We conduct the evolutionary search on the Minatar SpaceInvaders environment, which has previously been shown to facilitate the discovery of generalizable meta-RL algorithms (Jackson et al., 2024). For the search, we use GPT-4.1-mini, selected for its strong performance–cost tradeoff. Throughout evolution, we fix the PPO and discriminator hyperparameters, evaluating 200 candidate RA functions over a span of roughly three hours. Full details of the evolutionary hyperparameters are provided in Appendix D.
@@ 6.2 Generalization
We now turn to the out-of-distribution performance of DAIL. As a first step, we benchmark our method against prior RA functions (Table 1) on Brax and Minatar environments (excluding Minatar SpaceInvaders). All methods share identical hyperparameters, differing only in the choice of RA function. Aggregated scores (Agarwal et al., 2021) are shown in Figure 4. Among the baselines, AIRL performs slightly better than GAIL on Brax but—along with GAIL-heuristic—performs poorly on Minatar, likely because their reward assignment functions yield predominantly negative rewards, incentivizing agents to terminate episodes early. FAIRL performs poorly across both suites, likely due to its exponential, unbounded reward decay for positive log-ratios, which destabilizes training. Overall, the baseline trends align with observations from (Orsini et al., 2021).
Figure 4 demonstrates DAIL’s effectiveness across both benchmark suites. On Minatar, DAIL significantly outperforms all baselines across all evaluation metrics. On Brax, DAIL outperforms baselines on most metrics, with a slightly lower median than AIRL and statistically significant gains in mean performance. To quantify the robustness of these improvements on Brax, we employ the probability of improvement (PI) (Agarwal et al., 2021)—a metric that estimates the probability of DAIL outperforming a baseline on a randomly chosen task. The results in Figure 5 (left) show statistically significant PI values higher than $0.5$, providing further evidence of DAIL’s superiority over existing methods. To assess DAIL’s generalization beyond the policy optimizer (PPO) used during evolution, we evaluate its performance with A2C (Mnih et al., 2016). As shown in Figure 5 (right), DAIL maintains significant performance advantages over GAIL showcasing its effectiveness across different policy optimization algorithms.
@@ 7 Conclusion
While DAIL demonstrates strong generalization, the discovered RA function $r_{\text{disc}}$ does not correspond to a valid $f$-divergence and therefore lacks theoretical guarantees. Moreover, despite its strong performance, the RA function of DAIL remains static throughout training and does not adapt to the training state (e.g., number of updates remaining, loss, observed log-ratios). Exploring time-aware RA functions—those that condition on the training state—could yield richer, more informative learning signals, similar to (Jackson et al., 2024). Additionally, including more information into the LLM’s context such as environment information and training state may facilitate more effective crossovers.
@@ A.2 FAIRL Reward Assignment Function
Note: Prior works (Ho & Ermon, 2016; Ghasemipour et al., 2020; Fu et al., 2018; Garg et al., 2021) formulate the $f$-divergence minimization objective as $D_{f}(\rho_{\pi}\|\rho_{E})$, that is, the divergence of the learned policy’s occupancy measure $\rho_{\pi}$ from that of the expert $\rho_{E}$. Under this formulation, the FAIRL reward corresponds to the forward KL divergence.
@@ A.3 Reward Assignment Functions in GAIL and AIRL
which corresponds to the reward used in GAIL (Ho & Ermon, 2016).
@@ Appendix F Discovered Reward Assignment Functions
Table 6: Top 5 reward assignment functions discovered on MinAtar SpaceInvaders. Each function is expressed in terms of discriminator logits $l$.
Reward Assignment Functions
1. $\sigma(l)\cdot 0.5\cdot\big(\tanh(l)+1\big)$
2. $\min\Bigg(1.5,\max\Bigg(0,\,\begin{cases}0.5+0.8\,l-\frac{\mathrm{softplus}\big(1.5(-l-0.8)\big)}{1.5},&l\leq-0.8\\[8.0pt] 0.5+0.8\,l,&-0.8<l<0.8\\[8.0pt] 0.5+0.8\cdot 0.8+\frac{\mathrm{softplus}\big(1.5(l-0.8)\big)}{1.5},&l\geq 0.8\end{cases}\Bigg)\Bigg)$
3. $\text{softplus}(l)\cdot\sigma(1.5l)+0.5\cdot\mathrm{gelu}(l)$
4. $\frac{l}{1+|l|}\cdot\sigma(3l)\cdot 0.5\cdot\big(\tanh(l)+1\big)$
5. $0.5\cdot\bigg(\frac{l}{1+|l|}+1\bigg)\cdot\sigma(3l)$
@@ Appendix G Runtime and Compute Used
Table 7: Compute usage across environments. All reported times reflect wallclock time using 2 H100 NVL GPUs.
Environment Inner Loop (16 agents)
Wallclock time (s)
@@
Appendix H Comparison with non-adversarial methods
We additionally compare DAIL with non-adversarial methods, namely Behavior Cloning (BC) Pomerleau (1988) and IQ-Learn Garg et al. (2021), a state-of-the-art non-adversarial algorithm. For IQ-Learn, we use the official implementation provided by Garg et al. (2021). Results are reported in Table 8. DAIL consistently outperforms both BC and IQ-Learn. Interestingly, IQ-Learn fails to surpass BC in 3 of the 4 tested environments, a phenomenon also observed in prior work Lai et al. (2024); Jain et al. (2024), highlighting stability challenges inherent in non-adversarial imitation learning methods as well.
@@ Appendix H Comparison with non-adversarial methods
Method Ant HalfCheetah Hopper Walker2d
BC $23\pm 0.04 0.0$$09\pm 0.02$ $23\pm 0.09 0.0$$04\pm 0.01$
IQ-learn $32\pm 0.01 0.0$$2\pm 0.00 0.0$$2\pm 0.06 0.0$$03\pm 0.01$
DAIL $0.88\pm 0.06$ $0.90\pm 0.01$ $1.72\pm 0.04$ $0.97\pm 0.01$
Appendix I LLM Usage

### v2

@@ 6.1 Evolutionary Search
We performed the evolutionary search on the Minatar SpaceInvaders environment, which has previously been shown to facilitate the discovery of generalizable meta-RL algorithms (Jackson et al., 2024). For the search, we use GPT-4.1-mini, selected for its strong performance–cost tradeoff. Throughout evolution, we fix the PPO and discriminator hyperparameters, evaluating 200 candidate RA functions over a span of three hours. Full details of the evolutionary hyperparameters are provided in Appendix D.
@@ 6.2 Generalization
We now turn to the out-of-distribution performance of DAIL. As a first step, we benchmark our method against previous RA functions (Table 1) in Brax and Minatar environments (excluding Minatar SpaceInvaders). All methods share identical hyperparameters, differing only in the choice of RA function. Aggregated scores using the rliable library (Agarwal et al., 2021) are shown in Figure 4. Among the baselines, AIRL performs slightly better than GAIL on Brax but performs poorly on Minatar, similar to GAIL-heuristic. This is likely because their reward assignment functions yield predominantly negative rewards, incentivizing agents to terminate episodes early. FAIRL performs poorly across both suites, likely due to its exponential, unbounded reward decay for positive log-ratios, which destabilizes training. Overall, the baseline trends align with the observations from (Orsini et al., 2021).
Figure 4 demonstrates DAIL’s effectiveness across both benchmark suites. On Minatar, DAIL significantly outperforms all baselines across all evaluation metrics. On Brax, DAIL outperforms baselines on most metrics, with a slightly lower median than AIRL and statistically significant gains in mean performance. To quantify the robustness of these improvements on Brax, we employ the probability of improvement (Agarwal et al., 2021) metric, which estimates the probability that DAIL outperforms a baseline on a randomly chosen task. As shown in Figure 5 (left), DAIL achieves a probability of improvement greater than $0.5$ against all baselines, providing further evidence of its superior performance. To assess DAIL’s generalization beyond the policy optimizer (PPO) used during evolution, we evaluate its performance with A2C (Mnih et al., 2016). As shown in Figure 5 (right), DAIL maintains significant performance advantages over GAIL showcasing its effectiveness across different policy optimization algorithms. Finally, we assess DAIL’s generalization performance under various discriminator regularization strategies proposed by Orsini et al. (2021), including weight decay, an additional entropy bonus, and spectral normalization of the discriminator weights. We also consider the case without any regularization. As shown in Table 2, DAIL outperforms GAIL in 3 out of 5 regularization regimes.
@@
Algo Env none w-decay entropy spectral grad-pen DAIL Asterix 0.88 $\pm$ 0.03 1.33 $\pm$ 0.03 0.12 $\pm$ 0.01 0.92 $\pm$ 0.03 0.66 $\pm$ 0.03 Breakout 0.81 $\pm$ 0.07 0.74 $\pm$ 0.08 0.91 $\pm$ 0.02 0.77 $\pm$ 0.07 1.01 $\pm$ 0.00 SpaceInvaders 0.71 $\pm$ 0.07 0.81 $\pm$ 0.01 0.80 $\pm$ 0.01 0.70 $\pm$ 0.09 0.90 $\pm$ 0.00 Overall 0.80 $\pm$ 0.03 0.96 $\pm$ 0.03 0.61 $\pm$ 0.01 0.80 $\pm$ 0.04 0.85 $\pm$ 0.01 GAIL Asterix 1.18 $\pm$ 0.03 1.44 $\pm$ 0.03 0.48 $\pm$ 0.03 0.22 $\pm$ 0.03 0.52 $\pm$ 0.04 Breakout 0.76 $\pm$ 0.07 0.52 $\pm$ 0.10 0.89 $\pm$ 0.01 0.33 $\pm$ 0.10 0.85 $\pm$ 0.07 SpaceInvaders 0.61 $\pm$ 0.09 0.34 $\pm$ 0.09 0.81 $\pm$ 0.00 0.42 $\pm$ 0.08 0.81 $\pm$ 0.03 Overall 0.85 $\pm$ 0.04 0.76 $\pm$ 0.04 0.73 $\pm$ 0.01 0.32 $\pm$ 0.05 0.73 $\pm$ 0.03
Table 2: Performance of DAIL and GAIL under different discriminator regularization strategies. Hyperparameters are adopted from Orsini et al. (2021). Reported values denote the mean and standard error across runs.
@@ 6.3 Analysis
6.4 Stability
To assess the stability of the evolutionary process, we conduct an additional independent evolutionary run on the Minatar SpaceInvaders environment. The top-5 RA functions discovered in both runs are presented in Table 8 and the top-3 are plotted in Figure 8. We see that the $6$ evolved RA functions exhibit highly similar structures. Further, they maintain informative gradients within the range [1,0][-1,0] and saturate rapidly thereafter, consistent with our analysis in Section 6.3. Furthermore, we conduct an evolutionary run in the MinAtar Breakout environment and find that DAIL emerges among the top-5 RA functions in the final generation, demonstrating both the stability of the evolutionary process and the effectiveness of DAIL.
Figure 8: Comparing the top-33 discovered RA functions across two independant evolutionary runs. Note that Run 1 - #1 corresponds to DAIL.
@@ 7 Conclusion
While DAIL demonstrates strong generalization, the discovered RA function $r_{\text{disc}}$ does not correspond to a valid $f$-divergence and therefore lacks theoretical guarantees. Moreover, despite its strong performance, the RA function of DAIL remains static throughout training and does not adapt to the training state (e.g., number of updates remaining, loss, observed log-ratios). Exploring time-aware RA functions—those that condition on the training state—could yield richer, more informative learning signals, similar to Jackson et al. (2024). Additionally, including more information into the LLM’s context such as environment information may facilitate more effective crossovers.
Finally, it would be valuable to evaluate DAIL’s generalization to more complex benchmarks, such as the Atari-57 (Bellemare et al., 2013) and Procgen (Cobbe et al., 2020) suites. Prior work in meta-learning Oh et al. (2020) has demonstrated that training across a diverse set of environments yields more robust algorithms. For instance, Oh et al. (2025) introduced the Disco57 and Disco103 suites, consisting of 57 and 103 environments, respectively. Investigating the discovery of AIL algorithms meta-trained on such large-scale environment collections is an exciting direction for future research, though it currently lies beyond our computational budget.
@@ 9 Reproducibility statement
10 Acknowledgments
This research/project was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 24-SIS-SMU-010).
@@ A.3 Reward Assignment Functions in GAIL and AIRL
which corresponds to the reward used in GAIL (Ho and Ermon, 2016).
By plugging the $f$-functions of other commonly used $f$-divergences into Eq. 29, we can derive their corresponding reward assignment functions, as summarized in Table 3. While no formal adversarial imitation learning (AIL) methods explicitly employ these divergences, they have been explored in the context of non-adversarial imitation learning algorithms such as IQ-Learn (Garg et al., 2021).
Table 3: Reward assignment functions derived from different $f$-divergences
Divergence $f(x)$ $r_{f\text{-var}}$ Forward KL $\log x$ $\frac{\rho_{E}}{\rho_{\pi}}$ Reverse KL $x\log x$ $\log\frac{\rho_{E}}{\rho_{\pi}}-1$ Jensen-Shannon $x\log x-(x+1)\log\left(\frac{x+1}{2}\right)$ $\log\frac{1}{2}\left(1+\frac{\rho_{E}}{\rho_{\pi}}\right)$ Squared Hellinger $(\sqrt{x}-1)^{2}$ $\sqrt{\frac{\rho_{E}}{\rho_{\pi}}}-1$ Pearson $\chi^{2}$ $(x-1)^{2}$ $2\left(1-\frac{\rho_{\pi}}{\rho_{E}}\right)$ Total Variation $\frac{1}{2}|x-1|$ $\frac{1}{2}\cdot\operatorname{sign}\left(1-\frac{\rho_{\pi}}{\rho_{E}}\right)$
@@ Appendix F Discovered Reward Assignment Functions
Table 8: Top 5 reward assignment functions discovered generated after evolution. Each function is expressed in terms of discriminator logits $l$.
Environment Reward Assignment Functions SpaceInvaders-Run $1$ 1. $\sigma(l)\cdot 0.5\cdot\big(\tanh(l)+1\big)$ 2. $\min\Bigg(1.5,\max\Bigg(0,\,\begin{cases}0.5+0.8\,l-\frac{\mathrm{softplus}\big(1.5(-l-0.8)\big)}{1.5},&l\leq-0.8\\[8.0pt] 0.5+0.8\,l,&-0.8<l<0.8\\[8.0pt] 0.5+0.8\cdot 0.8+\frac{\mathrm{softplus}\big(1.5(l-0.8)\big)}{1.5},&l\geq 0.8\end{cases}\Bigg)\Bigg)$ 3. $\text{softplus}(l)\cdot\sigma(1.5l)+0.5\cdot\mathrm{gelu}(l)$ 4. $\frac{l}{1+|l|}\cdot\sigma(3l)\cdot 0.5\cdot\big(\tanh(l)+1\big)$ 5. $0.5\cdot\bigg(\frac{l}{1+|l|}+1\bigg)\cdot\sigma(3l)$ SpaceInvaders-Run 2 1. $\mathrm{clip}\!\left(0.5\cdot\big(\tanh(l)+1\big)\,\cdot\mathrm{softplus}(l)-0.1\,\mathrm{softplus}(-l),\,0,\,\infty\right)$ 2. $\mathrm{softplus}(l)\cdot\,\Big(1-\tanh^{2}\!\Big(\frac{l-3}{1.5}\Big)\Big)$ 3. $\sigma(3l)\cdot\,\mathrm{softplus}(l)$ 4. $\sigma\!\big(3(l-1)\big)\,\cdot\sigma\!\big(5(2.5-|l|)\big)$ 5. $\mathrm{clip}\Bigg(\big(\frac{l}{1+|l|}+0.2\!\left(\frac{l}{1+|l|}\right)^{\!3}+0.3\big)\cdot\exp\!\big(-0.2\,\max(l,0)\big),0,\infty\Bigg)$ Breakout 1. $0.5\cdot\big(\tanh(1.5\cdot l)+1\big)$ 2. $\mathrm{softplus}(l)\cdot\frac{l\cdot\sigma(l)+1}{2}$ 3. $\big(\tanh(l)+1\big)\cdot\sigma(l)$ 4. $\frac{\tanh(2.0\cdot l)}{2.0}+0.3\cdot\mathrm{softplus}(l)+0.5$ 5. $(1-\sigma(2.0\cdot l))\cdot 0.7\cdot\mathrm{softplus}(l)+\sigma(2.0\cdot l)\cdot\big(\mathrm{softplus}(l)+\mathrm{clip}(0.3\cdot l,-1,2)\big)$
@@ Appendix G Runtime and Compute Used
Table 9: All reported runtimes correspond to wall-clock time measured while training 16 agents in parallel on 2 H100 NVL GPUs.
Environment Wallclock time (s)
@@
Table 10: Comparison between DAIL and OPEN-ES across MinAtar and Brax environments. Reported values denote mean $\pm$ standard error over evaluation runs. Note that DAIL and OPEN-ES were meta-trained on MinAtar SpaceInvaders.
Environment DAIL OPEN-ES SpaceInvaders 0.90 $\pm$ 0.00 0.73 $\pm$ 0.05 Asterix 0.66 $\pm$ 0.03 1.27 $\pm$ 0.04 Breakout 1.01 $\pm$ 0.00 0.38 $\pm$ 0.10 HalfCheetah 0.90 $\pm$ 0.00 0.62 $\pm$ 0.02 Walker2d 0.95 $\pm$ 0.00 0.73 $\pm$ 0.02 Hopper 1.65 $\pm$ 0.01 1.09 $\pm$ 0.03 Reacher 0.63 $\pm$ 0.01 0.83 $\pm$ 0.01 Ant 0.85 $\pm$ 0.02 0.51 $\pm$ 0.03
Appendix H Comparison with Open-ES
Figure 10: RA function generated by OpenES
We evaluate the contribution of the LLM in guiding the evolutionary search by comparing it to a widely used baseline for optimizing the outer-loop objective (Eq.3): OpenAI Evolution Strategies (ES) Salimans et al. (2017). OpenAI-ES is a black-box, gradient-free optimization algorithm known for its strong performance on similar classes of problems Goldie et al. (2024); Sapora et al. (2024). For a fair comparison, both methods are given the same computational budget ($200$ inner-loop evaluations). From Table 10, we observe that OPEN-ES underperforms DAIL both on the training environment (SpaceInvaders) and on the test environments overall, consistent with the findings reported by Goldie et al. (2025). Figure 10 visualizes the discovered RA function, which appears irregular and non-smooth, likely contributing to its inferior performance.
Appendix I Comparison with additional divergence based methods
Table 11: Comparison with additional divergence-based AIL methods across MinAtar and Brax environments. Reported values denote mean $\pm$ standard error.
Environment DAIL Pearson Sq-Hellinger TV WGAIL Asterix 0.66 $\pm$ 0.03 -0.03 $\pm$ 0.00 -0.02 $\pm$ 0.00 -0.01 $\pm$ 0.00 0.52 $\pm$ 0.04 Breakout 1.01 $\pm$ 0.00 -0.01 $\pm$ 0.00 -0.01 $\pm$ 0.00 -0.01 $\pm$ 0.00 0.91 $\pm$ 0.06 Ant 0.85 $\pm$ 0.02 0.57 $\pm$ 0.02 0.82 $\pm$ 0.03 0.19 $\pm$ 0.05 0.80 $\pm$ 0.02 HalfCheetah 0.90 $\pm$ 0.00 0.49 $\pm$ 0.04 0.86 $\pm$ 0.01 -0.01 $\pm$ 0.01 0.87 $\pm$ 0.01 Hopper 1.65 $\pm$ 0.01 0.90 $\pm$ 0.11 1.49 $\pm$ 0.04 1.59 $\pm$ 0.04 1.50 $\pm$ 0.01 Reacher 0.63 $\pm$ 0.01 0.88 $\pm$ 0.05 0.94 $\pm$ 0.01 0.86 $\pm$ 0.07 0.42 $\pm$ 0.01 Walker2d 0.95 $\pm$ 0.00 0.54 $\pm$ 0.05 0.98 $\pm$ 0.00 0.91 $\pm$ 0.01 0.89 $\pm$ 0.01
We evaluate RA functions derived from alternative $f$-divergences, as summarized in Table 3, which has not been explored in prior work Ghasemipour et al. (2020); Orsini et al. (2021). We also compare DAIL to Wasserstein-GAIL (WGAIL) , which corresponds to replacing the discriminator objective (36) with:
$D^{*}(s,a)=\arg\max_{D\in\{f:\|f\|_{L}\leq 1\}}\;\mathbb{E}_{(s,a)\sim\rho_{E}}\left[D(s,a)\right]-\mathbb{E}_{(s,a)\sim\rho}\left[D(s,a)\right]$ (38)
Note that we already employ gradient penalty as a regularizer, enforcing the discriminator to be $1$-Lipschitz. After a hyperparameter sweep over the discriminator learning rate, we find that $3\times 10^{-3}$ performs best for WGAIL. From Table 11, we observe that DAIL outperforms the divergence-based AIL baselines on 5 out of 7 test environments.
Appendix J Comparison with non-adversarial methods
We additionally compare DAIL with non-adversarial methods, namely Behavior Cloning (BC;  Pomerleau (1988)) and IQ-Learn (Garg et al., 2021), a state-of-the-art non-adversarial algorithm. For IQ-Learn, we use the official implementation provided by Garg et al. (2021). Results are reported in Table 12. DAIL consistently outperforms both BC and IQ-Learn. Interestingly, IQ-Learn fails to surpass BC in 3 of the 4 tested environments, a phenomenon also observed in prior work (Lai et al., 2024; Jain et al., 2024), highlighting stability challenges inherent in non-adversarial imitation learning methods as well.
@@ Appendix I Comparison with additional divergence based methods
Method Ant HalfCheetah Hopper Walker2d BC $0.23$ $\pm$ 0.04 $0.09$ $\pm$ 0.02 $0.23$ $\pm$ 0.09 $0.04$ $\pm$ 0.01 IQ-learn 0.32-0.32 $\pm$ 0.01 0.02-0.02 $\pm$ 0.00 $0.02$ $\pm$ 0.06 $0.03$ $\pm$ 0.01 DAIL $0.88$ $\pm$ 0.06 $0.90$ $\pm$ 0.01 $1.72$ $\pm$ 0.04 $0.97$ $\pm$ 0.01
Appendix K LLM Usage